Instruction: Use the seven-step method to solve the following.

In a vegetables nursery, the caretaker thinks that the lengths

of seedlings in a box sprayed with a new kind of fertilization

has an average height of 16cm after three days and a

standard deviation of 5cm. One researcher randomly selected

60 such seedlings and calculated the mean height to be 12cm

and standard deviation was 5cm. Will you conduct a onetailed test or two tailed test. Proceed with the test using

α=0.05..


1
Expert's answer
2022-04-22T05:20:42-0400

H0 : "\mu"=16

H1 : "\mu"<16


we know that, mean = 12, n = 60 s = 5


z = (12-16)/(5/"\sqrt{60}") = -6.19

z alpha = -1.64

z < z alpha

So we reject H0 and accept H1

and we will conduct one tail test